18. Is the function f(x) = 6x + 2x + 7 even, odd, or neither?
Answer #1

Solution fex = 646+2x+7 of fun= f(x then Event Perde -fore, then odd other wise niether even her odd fl-H=61-4) 6+2(-x)+7 flo